There are 25 quarts of milk on a supermarket shelf, 6 of which are spoiled. A customer buys three quarts of milk.

Part A. How many samples contain at least one quart of spoiled milk?

Part B. How many samples contain no quarts of spoiled milk?
Answer #1

A) number of samples contain at least one quart of spoiled milk =total samples-sample which contain no quart of spoiled milk

=(25C3)-(19C3) =2300-969 =1331

B)

amples contain no quarts of spoiled milk =N(select all 3 from 19 good quarts) =(19C3)=969